OpenStudy (milo123):

Medal What is the height of a cone with a radius of 2 inches and volume of 6 cubic inches? Round your answer to the tenths place. 0.5 inch 1.4 inches 3.0 inches 4.0 inches

OpenStudy (milo123):

b?

OpenStudy (schoolgeek15):

1.43239 in Given radius and volume calculate the height, slant height, lateral surface area and total surface area. Given r, V find h, s, L, A h = (3 * v) / (πr2)

OpenStudy (schoolgeek15):

This is what someone else said, Hope it helps :)

OpenStudy (mayankdevnani):

You're correct :D Thumbs up

OpenStudy (milo123):

Thanks

OpenStudy (will.h):

\[V = \pi * r^2 *\frac{ h }{ 3 }\] Solving for h we will have the following equation \[h = 3 * \frac{ V }{ \pi *r^2 }\] \[h = 3 * \frac{ 6 }{ \pi * 2^2 }\] \[h = 1.43\]

OpenStudy (milo123):

sorry i can't medal everyone lol